# Lightweight Embedded FPGA Deployment of Learned Image Compression with Knowledge Distillation and Hybrid Quantization

### v3

@@ IV-D Ablation Study
In the next section, we study the contribution of each component of our architecture through ablation: since Student-160 shows the best RD efficiency while meeting the 25 FPS latency target, we will focus on this model from now on. We perform an ablation study where we independently drop from our architecture the knowledge distillation part, our GDN implementation, the mixed precision quantization, the quantization aware training, and the fully pipelined hardware configuration. For each ablated component, we assess the drop in RD efficiency, latency and resource utilization.
@@ IV-D1 Knowledge Distillation
To start with, we assess the effect of the loss terms in Eq. 1 on the RD-efficiency of the knowledge distillation process. To this end, we fine-tune the K128 model and train it with 1) $\mathcal{L}_{\text{latent}}$ and $\mathcal{L}_{\text{perc}}$ KD losses 2) with only $\mathcal{L}_{\text{latent}}$ 3) only $\mathcal{L}_{\text{perc}}$ loss and finally with neither loss. Fig. 12 shows that not using the latent space loss results in an 11.62% increase in BD-Rate and a trivial drop in PSNR as the latent space loss typically helps guide the compressed representation of the student model (the learned latent representation) to be closer to that of the teacher model. This encourages the student model to learn better compression by distilling high-level information from the teacher, improving the rate (bitstream size). Whereas The output space loss focuses on the quality of the reconstructed image. It ensures that the student’s output (reconstructed image) closely matches the teacher’s output, guiding it to produce higher fidelity reconstructions with lower computational cost. As such, the absence of the outputs space loss results in a 0.36 dB drop in PSNR. Removing the entirety of the KD learning scheme results in a considerable loss in performance with a 25% increase in bpp and a significant drop of approximately 1 dB in PSNR compared to only +14% bpp and -0.3 dB with KD as seen in Fig.15. This demonstrates the essential role of both latent space and output space knowledge distillation losses in enhancing the performance of the student model. The significant increases in BD-rate and decreases in PSNR observed in their absence underscore the importance of these losses in achieving efficient compression and high-quality reconstruction. These findings affirm that incorporating knowledge distillation is vital for bridging the performance gap between smaller models and their larger counterparts, ultimately leading to the desired RD and FPS targets.
Figure 12: BD-rate and BD-PSNR results for Latent Space and Output Space ablation showing a 0.6 PSNR drop and 25% rate increase when the KD protocol is not applied
@@ IV-D3 Quantization strategy
a) Mixed Precision vs Single Precision: As a first experiment, instead of allocating 32 rather than 8 bits to the GDN parameters to boost its precision (mixed precision), we allocate 8 bits to all layers, i.e. we implement a single precision quantization scheme.
@@
Tables. VII and VI highlight the difference in RD performance due to mixed precision quantization, assigning higher precision representation to the quantization error sensitive GDN activation layers results in consistent improvement in PSNR with an average of approximately +0.3 dB across all model permutations. Fig. 15 and 14 better highlight the advantage of using hybrid quantization, a change of approximately +0.3 dB PSNR and - 7% BPP can be observed, the margin of improvement is thinner for smaller student models, this is likely due to the model’s inability to properly capitalize on the higher precision representation because of its lower number of input and output channels per layer.
b) QAT vs. PTQ Quantization: Next, we replace Quantization Aware Tuning where quantization is followed by tuning for a simpler Post Training Quantization (PTQ) scheme where the model is quantized post-training without any further fine-tuning. Fig.13 shows the RD curves for both quantization schemes: PTQ entails a large RD efficiency drop that we measured in a 40% BD-rate and a -1 dB BD-PSNR penalties. This experiment illustrates the benefits of integrating the quantization step in the learning loop when a multi-loss function is minimized at training time.
TABLE VI: Single Precision KD Distillation Results on Kodak 720p Inputs
@@
TABLE VII: Mixed Precision KD Distillation Results on Kodak 720p Inputs
@@
TABLE VIII: Model latency as a function of the implementation configuration for a single 720p image.
Component Sequential Pipelined
FPS ↑ DSP Efficiency ↑ Avg Bandwidth FPS ↑ DSP Efficiency ↑ Avg Bandwidth
Encoder 6.57 48% 4.9 GB/s 15.6 85% 10.8 GB/s
Decoder 4.27 45% 4.2 GB/s 11.7 68% 7.7 GB/s
Both 2.58 59% 6.8 GB/s 8.09 88% 9.6 GB/s
TABLE IX: Model latency as a function of the implementation configuration for a single 720p image on the 160K sutdent.
Component Sequential Pipelined
FPS ↑ DSP Efficiency ↑ Avg Bandwidth FPS ↑ DSP Efficiency ↑ Avg Bandwidth
Encoder 10.1 56% 3.5 GB/s 43.6 87% 7.1 GB/s
Decoder 6.5 51% 3.3 GB/s 15.5 75% 6.7 GB/s
Both 3.8 59% 3.7 GB/s 22.1 81% 8.5 GB/s
@@ IV-D4 Sequential vs Pipelined:
Next, we assess the impact of our fully pipelined implementation in terms of latency over a simpler sequential configuration. We recall that our hardware design is fully pipelined as seen in Fig. 7, allowing parallel operations to occur simultaneously and taking full advantage of the three DPU cores. In contrast, other implementation in the literature in [20] use a unified computational lane that takes advantage of all available resources, leading to inefficiencies in bandwidth and resource utilization as a result of idle resources. Therefore, we experiment with a baseline sequential configuration where all model components (main encoder, hyperprior encoder, entropy encoder-decoder, hyperprior decoder, and main decoder) are executed one after another. This approach allows each component to utilize all available DPU cores for its specific task, yet at the cost of sub-optimal resource utilization due to idle resource usage within intermediate stages, leading to inefficiencies. Namely,a unified computational lane is dedicated to all processes, each module is processed individually and intermediate results are saved in the external memory and fetched when needed, all resources are available to whatever module is being run. Tables.VIII gives system level performance numbers for the mapped K192 teacher model using both design space configurations, the fully pipelined configuration results in higher DSP efficiency and external memory bandwidth which translate into a x2.5 FPS performance. In Table IX, the 160K student model, distilled via knowledge distillation, demonstrates significantly higher performance. KD reduces the model size, directly lowering the number of floating-point operations (FLOPs) required for inference. This reduction in computational load results in higher memory bandwidth (10.8 GB/s in the pipelined encoder) and improved DSP efficiency (87%) by allowing a more efficient utilization of available hardware resources on the ZCU102 FPGA. As a result, the student model achieves 42.6 FPS in the pipelined encoder, nearly 3x faster than the sequential model. The exploration of FPGA engine design reveals that the fully pipelined configuration offers substantial performance benefits compared to the sequential approach. By leveraging parallel processing and optimizing resource allocation, the pipelined setup fully exploits the capabilities of the ZCU102 FPGA.
@@ V Conclusion and future works
This work presents a hardware-efficient FPGA implementation of a Learned Image Compression (LIC) model, achieving real-time latency while maintaining high RD-efficiency. A novel transfer learning scheme is introduced, enabling the use of larger models while reducing computational requirements. Another contribution to the design is a custom Generalized Divisive Normalization (GDN) layer optimized for FPGA, significantly enhancing rate-distortion (RD) performance. A mixed-precision approach, using 32-bit precision for the GDN layer and 8-bit for the remaining layers, provides an optimal trade-off between resource usage and RD performance. This results in an average PSNR gain of 0.8 dB and a 7% reduction in bit-rate, with the GDN layer requiring just 4% of total model FLOPS. Additional optimizations through Quantization-Aware Training (QAT) and structured pruning further reduce model size and complexity without significantly impacting quality. QAT improves robustness to quantization, yielding a 40% BD-rate improvement and a 1 dB PSNR gain over standard Post-Training Quantization (PTQ). Structured pruning eliminates up to 30% of model parameters, enhancing resource efficiency and throughput on FPGA. These enhancements enable real-time performance, reaching 43 FPS on 720p inputs, while surpassing state-of-the-art RD metrics. Finally, our fully pipelined FPGA design maximizes resource utilization, delivering up to 3x the performance of sequential designs and closely matching floating-point reference models, thus outperforming other FPGA-based approaches in the literature. Future work will extend the knowledge distillation scheme to cross-architectural KD, leveraging the strong RD performance of vision transformers while distilling their knowledge into more efficient convolutional neural network autoencoder-based architectures. We also plan to refine the KD scheme to account for gradual dynamic weighting instead of discrete phases, and expand the tests to include the hyperprior model with the aim of improving the entropy performance of the student model. On the hardware side, we plan to explore low-level design space exploration using analytical and experimental software-hardware co-modeling to optimize mapping configurations.

### v5

@@ IV-D Ablation Study
In the next section, we perform an ablation study on the Student-160 model, which offers the best RD efficiency while meeting the 25 FPS latency target. We independently remove components including knowledge distillation, GDN implementation, mixed precision quantization, quantization-aware training, and the fully pipelined hardware configuration. For each ablation, we evaluate the impact on RD efficiency, latency, and resource utilization.
@@ IV-D1 Knowledge Distillation
We assess the impact of the KD loss terms in Eq. 1 on RD-efficiency by fine-tuning the K128 model with the following configurations: 1) both $\mathcal{L}_{\text{latent}}$ and $\mathcal{L}_{\text{perc}}$, 2) only $\mathcal{L}_{\text{latent}}$, 3) only $\mathcal{L}_{\text{perc}}$, and 4) no loss. Fig. 12 shows that omitting the latent space loss increases BD-Rate by 14.37% and causes a minor drop in PSNR, as the latent loss helps align the student’s learned latent space with the teacher’s, improving compression.
Whereas the output space loss ensures the student’s reconstructed image matches the teacher’s, improving fidelity with lower computational cost. Omitting this loss leads to a 0.54 dB PSNR drop. Removing the entire KD scheme results in a 22.44% increase in bpp and a 0.76 dB PSNR drop, compared to a +14% bpp and -0.3 dB drop with KD, as shown in Fig. 17. This highlights the critical role of both latent and output space KD losses in improving student model performance. The significant BD-rate increase and PSNR drop without these losses emphasize their importance for efficient compression and high-quality reconstruction. These results confirm that KD is essential for closing the performance gap between smaller and larger models, achieving the desired RD and FPS targets.
Figure 12: BD-rate and BD-PSNR results for Latent Space and Output Space ablation showing a 0.76 PSNR drop and 22.4% rate increase when the KD protocol is not applied
Now, we investigate the impact of fixing the teacher model during knowledge distillation using two pre-trained teachers from the CompressAI model zoo, differing in $\lambda$ values: one prioritizing image quality ($\lambda=0.0025$, red model in Fig.13) and the other minimizing bit rate ($\lambda=0.18$, red model in Fig.14). In previous experiments, we always use the same $\lambda$ for the Teacher-Student pair. The students, each with half the number of filters of the teachers, are trained with these two different teachers to analyze how the teacher’s focus affects the students’ RD performance. When the teacher emphasizes image quality, smaller student models benefit from higher PSNR than their pre-trained $\lambda$ counterparts, despite having fewer channels. However, this comes at the expense of increased bit rate. As $\lambda$ decreases, the effect becomes less pronounced, with the student models showing a +8.75% BD-rate and -0.39 dB BD-PSNR compared to the baseline teacher. In contrast, when the teacher focuses on minimizing bit rate, the students are pushed towards lower bit rates but suffer a reduction in image quality. This trade-off is more pronounced as $\lambda$ increases, resulting in a +10.32% BD-rate and -0.42 dB BD-PSNR compared to the baseline teacher.
While knowledge distillation benefits smaller models over training from scratch, its effectiveness depends on both the teacher’s initial performance, "priorities" and the student model size. Students trained with an distortion-focused teacher, seen in Fig.13 show improved PSNR but still cannot match the teacher’s RD performance due to the trade-off in rate. On the other hand, students trained with a rate-focused teacher, Fig.14, achieve lower bit rates at the cost of image quality. This experiment highlights that the benefits of KD are shaped by the teacher’s priorities and the student model’s size, particularly when balancing rate and distortion and the possibility of using a teacher-student pair with varying $\lambda$ values based on the objective.
Figure 13: BD-rate and BD-PSNR results
Figure 14: BD-rate and BD-PSNR results
@@ IV-D3 Quantization strategy
a) Mixed Precision vs Single Precision: As a first experiment, instead of allocating 32 rather than 8 bits to the GDN parameters to boost its precision (mixed precision), we allocate 8 bits to all layers, i.e. we implement a single precision quantization scheme. Assigning higher precision to the quantization-sensitive GDN activation layers consistently improves PSNR, averaging +0.3 dB across all model configurations. Fig. 17 and 16 illustrate the benefits of hybrid quantization, with a +0.2 dB PSNR improvement and -2% BPP. The improvement is less significant for smaller student models, likely due to their limited ability to leverage the higher precision representation, given their fewer input and output channels per layer.
@@
b) QAT vs. PTQ Quantization: Next, we replace Quantization Aware Training where quantization is jointly carried out with a fine-tuning step for a simpler Post Training Quantization (PTQ) scheme where the model is quantized post-training without any further fine-tuning. Fig. 15 shows the RD curves for both quantization schemes. Compared to QAT, PTQ results in a significant RD efficiency drop, with a 40% increase in BD-rate and a -1 dB decrease in BD-PSNR. This experiment highlights the advantages of integrating the quantization step into the learning loop when minimizing a multi-loss function during training.
TABLE VIII: Single Precision KD Distillation Results on Kodak 720p Inputs
@@
TABLE IX: Mixed Precision KD Distillation Results on Kodak 720p Inputs
@@ IV-D4 Sequential vs Pipelined:
Next, we assess the impact of our fully pipelined implementation on latency compared to a simpler sequential configuration. Our design, shown in Fig. 7, enables parallel operations that fully exploit the three DPU cores. In contrast, other implementations in the literature [26] use a unified computational lane, leading to bandwidth and resource inefficiencies due to idle resources. We therefore test a baseline sequential configuration where all components (main encoder, hyperprior encoder, entropy encoder-decoder, hyperprior decoder, and main decoder) are executed consecutively. Although each component utilizes all DPU cores, this approach results in suboptimal resource usage during intermediate stages. A unified computational lane is dedicated to all processes, with each module processed sequentially and intermediate results stored in external memory as needed. Table X presents the performance of the K60 teacher model under both configurations. The pipelined design achieves 2.5× higher FPS thanks to improved DSP efficiency and external memory bandwidth. The distilled 160K student model shows significant gains, with knowledge distillation reducing model size and FLOPs, yielding 10.8 GB/s bandwidth and 87% DSP efficiency. The student model reaches 42.6 FPS in the pipelined encoder—nearly 3× faster than the sequential version. Overall, the pipelined design fully exploits the ZCU102 FPGA, offering substantial throughput benefits.
TABLE X: Model latency as a function of the implementation configuration for a single 720p image on the 160K student.
Sequential Pipelined
[…]
Encoder 17.3 56% 3.5 GB/s 43.6 87% 7.1 GB/s
Decoder 13.6 51% 3.3 GB/s 30.5 85% 6.7 GB/s
@@ V Conclusion and future works
This work presents a hardware-efficient FPGA implementation of a Learned Image Compression (LIC) model, achieving real-time latency and high RD efficiency. A novel Knowledge Distillation-based transfer learning scheme reduces computational cost while maintaining performance, using larger models as teachers to guide smaller models in jointly learning better reconstruction and rate representations. We introduce an FPGA-optimized Generalized Divisive Normalization (GDN) layer with a mixed-precision approach, yielding PSNR gains and bit-rate reduction while minimizing FLOPS. Optimizations through Quantization-Aware Training (QAT) and structured pruning reduce complexity, enhancing FPGA efficiency without sacrificing quality. The design employs patch-parallel pipelining, achieving real-time 720p performance and outperforming existing RD metrics with up to three times the performance of sequential designs. Future work will extend knowledge distillation by exploring alternative loss functions, including one for activation maps to better guide the student. We will also investigate freezing the hyperprior model, extending KD to the hyperprior, or focusing on distilling the decoder for edge applications. Further refinements in dynamic weighting and continued software-hardware co-modeling will optimize mapping configurations.